OpenStudy (anonymous):

why is this true? \[\cos(2\times2t)=2\cos ^{2}2t-1\]

OpenStudy (phi):

cos(a+b)= cos(a)cos(b) - sin(a) sin(b) if a and b are the same thing i.e. b= a then this is cos(a+a)= cos^2(a) - sin^2(a) also, you might know that sin^2(a) + cos^(a) =1 you can use this to replace the sin^2(a) in the above equation

OpenStudy (anonymous):

where did this come from? \[\cos (a+b)\]

OpenStudy (anonymous):

You might think of cos(a+b) =cos(a+a) when b=a. So, cos(a+a)=cos(2a) and you can apply the formula for cos(a+b)=cosacosb-sinasinb and then replace all your b's with a's. Once you have that, replace your a with 2t and you have eyour answer.

OpenStudy (phi):

the cos of the sum of 2 angles cos(a+b) is well-known and can be proven many ways. see http://www.proofwiki.org/wiki/Cosine_of_Sum

OpenStudy (anonymous):

ok, that is not what my confusing is, my original LHS was cos (2 times 2t), how did that turned into cos (2+2t)?

OpenStudy (phi):

2* x is the same as x+x in your case 2*2t is the same as 2t + 2t if we call 2t = a we can use cos(a+a)

OpenStudy (anonymous):

ok i got you, this is good new information, thanks!

OpenStudy (phi):

can you finish ?

OpenStudy (anonymous):

i take a look and let you know in a minute...

OpenStudy (phi):

in case the typo confuses you also, you might know that \[ \sin^2(a) + \cos^2(a) =1 \] you can use this to replace the sin^2(a) in the above equation

OpenStudy (anonymous):

ok i got it, i see, now i just have to work out the rest of the problem which is a verification problem, this was just one of its step, i might need you help, let you know, thanks so far!

OpenStudy (anonymous):

ok, im completely lost in all the formulas

OpenStudy (anonymous):

original question was \[\cos 4t = 8\cos ^{4}t-8\cos ^{2}t+1\]

OpenStudy (anonymous):

i got as far as you explained

OpenStudy (anonymous):

now im stuck at \[2 \cos ^{2}2t-1\]

OpenStudy (anonymous):

That's your left hand side. Your right hand side is in terms of powers of cosines, so see if you can get them all together and simplify it, perhaps look at factoring. That is usually the case when we are trying to solve a trigonometric equation.

OpenStudy (anonymous):

i need to verify that both side equal, according to the answer the RHS doesn't change

OpenStudy (anonymous):

Oh, forgot you are trying to verify an identity.

OpenStudy (anonymous):

In that case, I would have started with the more complicated side, which is the RHS.

OpenStudy (anonymous):

The right hand side factors, btw.

OpenStudy (anonymous):

ok im listening to your idea

OpenStudy (anonymous):

hang on, lemme double check my last statement!

OpenStudy (anonymous):

ok, that 8cos^4t is what is presenting an issue....I am looking at a few possible approaches....give me just a few more moments..don't want to lead you astray. But I usually work with the more complex side first.

OpenStudy (anonymous):

ok, take your time

OpenStudy (anonymous):

nvrmind...phi was correct in the approach.

OpenStudy (phi):

you started with \[\cos 4t = 8\cos ^{4}t-8\cos ^{2}t+1 \] as you know, we can replace cos(4t) with \[ 2 \cos ^{2}2t-1 \] let's focus on cos(2t). using the same formula \[ \cos(2t)=2\cos ^{2}t-1 \] so we can replaces cos(2t) with that mess to get \[ 2 \cos ^{2}2t-1 \text { becomes } 2(2\cos ^{2}t-1)^2 -1 \]

OpenStudy (phi):

can you multiply out \[ (2 \cos^2 t -1)^2 = (2 \cos^2 t -1)(2 \cos^2 t -1)\] ?

OpenStudy (anonymous):

im looking

OpenStudy (phi):

you can use FOIL (if you know that way) it is the same as (x-y)(x-y)= x^2 -2 y + y^2 except x is 2 cos^2(t) and y is 1

OpenStudy (anonymous):

this is crazy...

OpenStudy (anonymous):

almost got it! wow, this is crazy! thank you both for your help! i do hope this question won;t be on my test on tuesday...

OpenStudy (anonymous):

lol...anything is possible....not trying to scare you, btw..just need to be prepared for anything.

OpenStudy (phi):

Here is how to do it using the distributive property if the problem were (x+y)*A we would do x*A + y*A let's call A= (2 cos^2t -1) and multiply A times (2 cos^2t -1) (2 cos^2t -1)A= 2 cos^2t * A - A but A is (2 cos^2t -1), so this is \[ (2 \cos^2t) (2 \cos^2t-1) - (2 \cos^2t -1)\] now distribute again \[ 2\cos^2t \cdot 2 \cos^2t - 2 \cos^2t - 2 \cos^2t -1 \] or \[ 4 \cos^4t - 4 \cos^2t -1 \]

OpenStudy (anonymous):

Forgot the --1 at the end.

OpenStudy (phi):

yes calmat noticed that should be +1 in \[ 4 \cos^4t - 4 \cos^2t +1 \]

OpenStudy (phi):

simplify \[ 2 ( 4 \cos^4t - 4 \cos^2t +1) -1 \]

OpenStudy (anonymous):

great explanation! Thank you both again!

OpenStudy (phi):

to do this problem you had to use the cos(2x) --> 2 cos^2(x) -1 twice, to change cos(4x) into cos(x) (with exponents)